OpenStudy (anonymous):

Let f(x) = ax + b where a and b are real numbers, f(f(f(1))) = 29 and f(f(f(0))) = 2. Then b equals ?? A start would be good.

OpenStudy (anonymous):

help yourself xD

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

Sarah, if you're not going to help, at least don't insult the person asking.

OpenStudy (anonymous):

umm. i just realised.. shouldnt i just find f(f(f(x))) ??

OpenStudy (anonymous):

Haha. Its ok Trash.

OpenStudy (anonymous):

trashman its none of your business, he knows im joking so be quiet and dont you dare tell me what to do..

OpenStudy (anonymous):

f(f(f(x)))=a^3(x) + a^2b + b .. is that right?

OpenStudy (jdoe0001):

I have \(\large a^3+a^2b+ab+b \implies 29\)

OpenStudy (anonymous):

Oh yeah. I'm sorry, i messed it up. ^^

OpenStudy (anonymous):

Ok this should be easy now, ig? making factors and dividing the two equations? is that where we are going?

OpenStudy (jdoe0001):

I'd think so, yes

OpenStudy (anonymous):

Its not working out like that.. @jdoe0001

OpenStudy (jdoe0001):

a sec lemme check about

OpenStudy (anonymous):

That's a lot of f's o-O

OpenStudy (anonymous):

Yeah its like. f(f(x))= a (ax+b) + b.. you just keep putting the value of f(x) in the x..

OpenStudy (anonymous):

That sounds very confusing.. although f(x) is just a fancy way of saying y

OpenStudy (jdoe0001):

@ohyeahh what did you get for f(f(f(0))) ?

OpenStudy (anonymous):

a^2.b + ab +b = 2 .. thats right?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

now gimme a sec to paste, tis a system of equations, so you know :)

OpenStudy (anonymous):

Take your time. Thank you.

OpenStudy (anonymous):

f(x)=ax+b f(1)=a+b f(f(1))=f(a+b)=a(a+b)+b=a^2+ab+b f(f(f(1)))=a(a^2+ab+b)+b=a^3+a^2b+ab+b=29 f(0)=0*a+b=b f(f(0))=f(b)=ab+b f(f(f(0)))=f(ab+b)=a(ab+b)+b=a^2b+ab+b=2 a^3+a^2b+ab+b=29 a^2b+ab+b=2 a^3+2=29 a^3=29-2=27 a=3 (3^2+3+1)b=2 b=2/13

OpenStudy (jdoe0001):

$$ a^3+a^2b+ab+b=29\\ 0+a^2b+ab+b=2 \ \ (\times \color{red}{-1})\\ --------------\\ a^3+a^2b+ab+b=29\\ 0\color{red}{-a^2b-ab-b}=-2\\ ---------------\\ a^3=27 \implies a=\sqrt[3]{27} = 3 $$

OpenStudy (anonymous):

I guess finding a and then finding b should be easier? I really couldnt follow how you found b. I got confused by the time i reached the end.

OpenStudy (jdoe0001):

and from there, as you can see @surjithayer lines, you substitute (a) to get (b)

OpenStudy (anonymous):

lol sorry.. it was you who wrote that. it was surjith.. but your method looks easier. i mean i couldnt really follow him..

OpenStudy (anonymous):

not you* ughh .